OpenStudy (anonymous):

in square root of (2-x) minus square root of (2x) = square root of (2+x), does the final answer have roots?

OpenStudy (phi):

do you mean \[ \sqrt{2-x} - \sqrt{2x} = \sqrt{2+x}\]

OpenStudy (phi):

there is an integer answer for x